OpenStudy (anonymous):

Use your grapher to determine which of the graphs matches the polar equation r = 4 cos 3θ. I graphed it and mine didn't match any of the correct graphs. Could somebody help please?

OpenStudy (anonymous):

@shrinifores

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

@Mimi_x3 @ParthKohli

OpenStudy (anonymous):

i think this is the one.

OpenStudy (anonymous):

and by the way, x axis is (theta)

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

here are the graphs i'm given.

OpenStudy (anonymous):

None of my graphs looked anything like these. I don't get which is the correct answer

OpenStudy (unklerhaukus):

what image are you getting when using your grapher ?

OpenStudy (unklerhaukus):

have you set it to polar coordinates?

OpenStudy (anonymous):

this is what i got

OpenStudy (unklerhaukus):

that is r= cos(3)θ you want to graph r= cos(3θ)

OpenStudy (anonymous):

ohhh! Thank you so much!!!

OpenStudy (anonymous):

Could you help me with one more?

OpenStudy (unklerhaukus):

sure. ps, i like the shape of r= cos(3)θ better

OpenStudy (anonymous):

Do i post it on here or make a new question? & haha me too

OpenStudy (anonymous):

Find all polar coordinates of point P where P = (9, pi/5)

OpenStudy (unklerhaukus):